OpenStudy (yamyam70):

Factor out the GCF 4-4 ^n +1 should be 4 (1-4^n) but I'm thinking the answer is 4 (1-1^n) someone explain ?

OpenStudy (neer2890):

is it \[4-4^{n}+1\]??

OpenStudy (yamyam70):



OpenStudy (neer2890):

\[4-4^{n+1}=4-4*4^{n}=4(1-4^{n)}\]

OpenStudy (yamyam70):

thank you man!

OpenStudy (neer2890):

It's a property used here which states that whenever base is same powers are added. So, \[4^{1}*4^{n}=4^{n+1}\]

OpenStudy (neer2890):

i think you already know it.. you're welcome..!

OpenStudy (yamyam70):

yeah just a confusion thanks again :)

OpenStudy (neer2890):

nvm... just remember in future...:)

OpenStudy (yamyam70):

I will :)